40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 16 19:48:52 2016, Lydia, Update, General, Alignment status 
    Reply  Sat Sep 17 02:00:23 2016, rana, Update, SUS, Alignment status 
       Reply  Sat Sep 17 16:51:01 2016, Lydia, Update, SUS, Alignment status Screenshot_from_2016-09-17_16-45-00.png
          Reply  Sun Sep 18 16:18:05 2016, rana, Update, SUS, Alignment status 
             Reply  Mon Sep 19 11:11:43 2016, ericq, Update, SUS, Alignment status 
                Reply  Mon Sep 19 22:03:10 2016, ericq, Update, General, Further recovery progress 
                   Reply  Tue Sep 20 10:45:06 2016, rana, Update, General, Further recovery progress 
                      Reply  Wed Sep 21 01:08:02 2016, ericq, Update, General, Further recovery progress 
                         Reply  Wed Sep 21 09:04:57 2016, Steve, Update, General, 8 hours recovery progress 8hrs.png
                            Reply  Thu Sep 22 15:44:21 2016, Steve, Update, General, ITMX magnets are stucked again StuckAgain.png
                               Reply  Thu Sep 22 20:01:47 2016, Lydia, Update, General, ITMX freed, all optics kicked 
                                  Reply  Thu Sep 22 22:52:08 2016, ericq, Update, General, Damping found to be on 
Message ID: 12507     Entry time: Mon Sep 19 22:03:10 2016     In reply to: 12504     Reply to this: 12508
Author: ericq 
Type: Update 
Category: General 
Subject: Further recovery progress 

[ericq, Lydia, Teng]

Brief summary of this afternoon's activities:

  • PMC alignment adjusted (Transmission of 0.74)
  • IMC locked, hand aligned. Tranmission slightly over 15k. Measured spot positions to be all under 2mm.
  • Set DC offsets of MC2 Trans + WFS1 + WFS2 (WFS2 DC offsets had wandered so much that DC "centered" left some quadrants almost totally dark)
  • Set demod offsets of WFS1+WFS2
    • Note to self: WFS script area is a mess. I can never remember which scripts are the right ones to run. I should clean this up
  • WFS loops activated, tested. All clear.
  • Locked Yarm, dither aligned. Transmission 0.8
  • Moved BS to center ITMY reflection on AS camera
  • Misaligned ETMY, aligned PRM to make a flashing PRY AS beam. REFL camera spot confirmed to be on the screen, which is nice
  • Wandered ITMX around until its AS spot was found. ITMX OSEMs not too far from their half max. (todo: update with numbers)
  • Wandered SRM around until full DRMI flashes seen
  • Centered all vertex oplevs
  • Made a brief attempt at locking X arm, could only get some crazy high order mode to lock. BS and ITMX alignments have changed substantially from the in-air locks, so probably need to adjust ETMX much more.

Addendum: I had a suspicion that the alignment had moved so much, we were missing the TRX PDs. I misaligned the Y arm, and used AS110 as a proxy for X arm power, as we've done in the past for this kind of thing. Indeed, I could maximize the signal and lock a TM00 mode. Both the high gain PD and QPD in the TRX path are totally dark. This needs realignment on the end table.

ELOG V3.1.3-